# ASCoT: An Adaptive Self-Correction Chain-of-Thought Method for Late-Stage Fragility in LLMs

### v3

@@ A.3. Analysis of Late‑Stage Fragility
Figure 7. Example generations by ASCoT, using Qwen2.5-Instruct 7B.
A.4. Generation Examples
To provide a qualitative illustration of our method’s end-to-end functionality, we present representative examples of reasoning chains after processing by the ASCoT pipeline. The Figure 7 displays the corrected output after being processed by ASCoT. This transformation from a flawed to a correct reasoning chain highlights the core capabilities of our system. It demonstrates the AVM module successfully identifying a high-risk step by analyzing its content quality and positional impact, followed by the MSCE module executing a targeted repair. This example visually confirm ASCoT’s ability to structure reasoning but, more importantly, to dynamically detect and rectify internal inconsistencies, ultimately leading to a more trustworthy final answer.
A.5. Future Work

### v4

@@ E.2 Distribution of Results
Figure 7: Example generations by ASCoT, using Qwen2.5-Instruct 7B.
Appendix F Detailed Ablation Studies
In this section, we present a series of ablation experiments to validate the design choices of ASCoT. All experiments were conducted using the LLaMA-3.1-8B model on the GSM8K benchmark.
F.1 Impact of Core Components
We first investigate the necessity of the two fundamental stages: semantic pruning (IRM) and the correction pipeline (AVM+MSCE). Table 6 summarizes the results.
Table 6: Ablation of core components. “w/o IRM” denotes removing the pruning stage; w/o Correction denotes using only pruning without any verification or correction.
Configuration Acc. (%) Tokens ActRatio
ASCoT (Full) 86.9 214.24 1.00
w/o IRM 87.1 325.40 1.52
w/o Correction 82.5 198.12 0.92
As shown in Table 6, removing the IRM leads to a massive increase in token usage (+52%) with negligible accuracy gain (+0.2%), validating IRM’s critical role in efficiency. Conversely, removing the correction pipeline (w/o Correction) causes accuracy to drop significantly to 82.5%, underscoring that semantic pruning alone induces fragility that must be compensated for by our robust verification mechanism.
F.2 Efficacy of Position-Aware Verification
A key innovation of ASCoT is the integration of the Positional Impact Score within the AVM module to address Late-Stage Fragility. Table 7 compares our strategy against a uniform risk assessment.
Table 7: Ablation of the AVM strategy. We compare our position-aware scoring against a uniform weighting scheme.
AVM Strategy Acc. (%) $\Delta$ Acc.
Position-Aware (Ours) 86.9
Uniform Weights 84.3 -2.6
The results in Table 7 reveal a 2.6% performance degradation when position awareness is removed. This empirical evidence supports our hypothesis that errors in later reasoning steps are more detrimental. By assigning higher risk weights to these terminal steps, ASCoT effectively reallocates verification resources to where they are most needed.
F.3 Analysis of Dual-Path Correction
Finally, we dissect the Multi-Perspective Self-Correction Engine (MSCE) to evaluate the contribution of its Dual-Path (Intrinsic + Extrinsic) strategy compared to single-source baselines.
Table 8: Ablation of the MSCE correction strategy. “Intrinsic” relies on self-reflection; “Extrinsic” uses independent regeneration.
Correction Strategy Acc. (%) $\Delta$ Acc.
Dual-Path (Ours) 86.9
Extrinsic Only 86.1 -0.8
Intrinsic Only 85.0 -1.9
Table 8 demonstrates that the Dual-Path strategy outperforms individual approaches. Relying solely on Intrinsic correction yields the lowest accuracy (85.0%), likely because the model struggles to escape its own erroneous reasoning context. Extrinsic correction performs better (86.1%) but is surpassed by the synergistic combination of both, which leverages the strengths of context-aware reflection and objective regeneration.
Appendix G Generation Examples
To provide a qualitative illustration of our method’s end-to-end functionality, we present representative examples of reasoning chains after processing by the ASCoT pipeline. Figure 7 displays the corrected output after being processed by ASCoT. This transformation from a flawed to a correct reasoning chain highlights the core capabilities of our system. It demonstrates the AVM module successfully identifying a high risk step by analyzing its content quality and positional impact, followed by the MSCE module executing a targeted repair. This example confirms ASCoT’s ability to structure reasoning but, more importantly, to dynamically detect and rectify internal inconsistencies, ultimately leading to a more trustworthy final answer.
Appendix H Future Work